# USD: Unsupervised Soft Contrastive Learning for Fault Detection in Multivariate Time Series

### v1

@@ Abstract
Unsupervised fault detection in multivariate time series is critical for maintaining the integrity and efficiency of complex systems, with current methodologies largely focusing on statistical and machine learning techniques. However, these approaches often rest on the assumption that data distributions conform to Gaussian models, overlooking the diversity of patterns that can manifest in both normal and abnormal states, thereby diminishing discriminative performance. Our innovation addresses this limitation by introducing a combination of data augmentation and soft contrastive learning, specifically designed to capture the multifaceted nature of state behaviors more accurately. The data augmentation process enriches the dataset with varied representations of normal states, while soft contrastive learning fine-tunes the model’s sensitivity to the subtle differences between normal and abnormal patterns, enabling it to recognize a broader spectrum of anomalies. This dual strategy significantly boosts the model’s ability to distinguish between normal and abnormal states, leading to a marked improvement in fault detection performance across multiple datasets and settings, thereby setting a new benchmark for unsupervised fault detection in complex systems. The code of our method is available at https://github.com/zangzelin/code_USD.git.
@@ I Introduction
In the realm of industrial operations and complex systems management [25], unsupervised fault detection [33] in multivariate time series data holds paramount importance. This technology underpins the early identification of potential failures and anomalies, safeguarding against costly downtimes and ensuring the continuity of critical processes [15]. Current methodologies in this field span a broad spectrum, from traditional statistical approaches to advanced machine learning models, each aiming to autonomously detect irregularities without the need for labeled data [32, 27]. These techniques are crucial in environments where supervisory labels are scarce or infeasible to obtain, making unsupervised methods invaluable for real-time monitoring and preventive maintenance strategies [35].
However, despite the use of deep learning [18] mathods advanced unsupervised fault detection techniques, false alarms and missed alarms are still common in practice (as shown in Table. II). The main problem stems from the unsupervised approach itself, which avoids the laborious task of manually collecting labels, but also loses the ability to precisely define different states. Without explicit labels, the system must rely on certain assumptions to distinguish between normal and abnormal states. As a result, as shown in Fig. 1(a), many existing methods use the Gaussian assumption that abnormal states are outliers in a Gaussian distribution. While this simplifies the modeling process, it fails to capture the complexity and diversity of the actual operating states, which weakens the ability to accurately detect subtle differences between states and reduces the effectiveness of fault detection models.
Theoretically, as shown in Fig. 1(b), the Gaussian assumption ignores the fact that both normal and abnormal states may contain multiple sub-states. This lead to unnatural distortions in the shape of the data stream, which ultimately affects the performance of these methods. Practically speaking, the Gaussian assumption simplifies the modeling process but does not take into account the inherent complexity and diversity of patterns that characterize the operating state. This simplification reduces the discriminative power of fault detection models, masking subtle differences and transitions between normal operation and potential faults.
Figure 1: Motivation of proposed USD (this paper). (a) Baseline fault detection methods use Gaussian assumptions and are unable to accurately model fault detection latent spaces with multiple sub-states. (b) proposed idea of this paper, the proposed soft contrast learning is introduced for more accurate modeling to achieve more precise identification accuracy.
To address the above issues, we introduce an unsupervised soft contrast learning framework, named Unsupervised Soft Fault Detection (USD), to go beyond the limitations imposed by the Gaussian distribution assumption by using more accurate multi-Gaussian representations [34]. As shown in Fig. 1(b), we first design a neighbor-based unsupervised data augmentation method to generate positive samples for unsupervised contrast learning [16] through online data augmentation of multivariate time series data. Then, we compute the loss function for soft contrast learning by incorporating a representation learning head module. Soft contrast learning optimizes the discriminative representation space based on the similarity of local streamforms to better portray the difference between normal and abnormal states. We argue that incorporating soft contrast learning not only enhances the model’s ability to distinguish between normal and abnormal patterns, but also captures a wider range of potential anomalies [37]; thus providing a more accurate, robust and adaptive solution.
Experimental evaluations show that our model performance (AUC and PR) is significantly improved compared to existing fault detection frameworks. Specifically, our approach achieves new enhancements on multiple standard test benchmarks (SWaT [9], WADI [2]) and operational scenarios with lower false alarm rates and higher detection accuracies. At the same time, we demonstrate, e.g., through visualization, the plausibility of our multi-Gaussian assumptions and the fact that soft-contrast learning can characterize different sub-states well. We summarize the contributions of this paper as follows.
  • Proposes a more plausible assumption of multiple manifold. In this paper, by challenging the traditional reliance on Gaussian distribution assumptions, our approach recognizes and accommodates the existence of many different states under normal and abnormal operating conditions. This realization of the complexity of the state space is critical to the development of more detailed and effective fault detection systems.
  • Introduces a soft contrast learning framework for performing latent space representations. We introduce a novel combination of data augmentation and soft contrastive learning techniques to represent these diverse states accurately. This methodology not only enhances the precision of state representations but also ensures that the subtle differences and transitions between normal and abnormal conditions are captured more effectively.
  • Obtain a significant performance boost. The practical efficacy of our approach is demonstrated through significant performance improvements across multiple datasets, with our model outperforming existing benchmarks by over 5%. These advancements underscore our contributions towards developing a more adaptive, accurate, and robust framework for unsupervised fault detection, paving the way for future innovations in monitoring and maintaining complex systems.
@@ II-A Time Series Anomaly Detection
In the evolving landscape of time series anomaly detection, the reliance on One-Class Classification (OCC) based methods, such as those presented by Chalapathy and Chawla [5], Hundman et al. [13], and further exemplified by DeepSVDD [22], EncDecAD [20], OmniAnomaly [24], USAD [4], and DAEMON [7], has underscored a critical limitation in their operational premise. These methods, by design, necessitate a training dataset composed entirely of normal instances, a requirement that is not only challenging but often impractical in real-world scenarios. The presence of even a small fraction of anomalous data within the training set can significantly impair the model’s detection capabilities, as highlighted by Wang et al. [26] and Huyan et al. [14]. This recognition of inherent limitations has spurred the exploration of alternative approaches that can circumvent the constraints imposed by OCC-based frameworks.
Responding to this challenge, Dai and Chen introduced GANF [8], a pioneering method that adopts an unsupervised learning strategy for the detection of Multivariate Time Series (MTS) anomalies. Qihang Zhou introduced MTGFlow [34], an unsupervised anomaly detection approach for Multivariate Time Series, leveraging dynamic graph structures and entity-aware normalizing flows for density estimation. The method notably incorporates MTGFlow-cluster [35], a strategy that clusters entities with similar characteristics to enhance precision, achieving superior performance on benchmark datasets. Chen Liu [19] proposed AnomalyLLM, a knowledge distillation-based time series anomaly detection method utilizing a large language model (LLM) as a teacher to guide a student network. Incorporating prototypical signals and synthetic anomalies, AnomalyLLM significantly improves detection accuracy by at least 14.5% on the UCR dataset, showcasing its effectiveness across 15 datasets.
@@ II-B Contrastive Learning and Soft Contrastive Learning
Soft contrastive learning [31, 29] and deep manifold learning [21, 17] represent two advanced methodologies in the domain of machine learning, particularly in the tasks of unsupervised learning and representation learning. These techniques aim to leverage the intrinsic structure of the data to learn meaningful and discriminative features.
At its core, soft contrastive learning [31, 29] is an extension of the contrastive learning framework [6, 11, 12], which aims to learn representations by bringing similar samples closer and pushing dissimilar samples apart in the representation space. However, soft contrastive learning introduces a more nuanced approach by incorporating the degrees of similarity between samples into the learning process, rather than treating similarity as a binary concept. This is achieved through the use of soft labels or continuous similarity scores, allowing the model to learn richer and more flexible representations. The softness in the approach accounts for the varying degrees of relevance or similarity among data points, making it particularly useful in tasks where the relationship between samples is not strictly binary or categorical, such as in semi-supervised learning or in scenarios with noisy labels.
@@ III Preliminary
In this section, we give a brief introduction of normalizing flow based false detection to better understand our proposed USD.
Figure 2: Overview of the proposed USD. (a) The framework of USD. First, USD samples $x$ captured within a sliding window of size $T$ along with their augmented positive counterparts $x^{+}$ are fed into the RNN $R(\cdot)$ to capture the temporal correlations within the data, generating the temporal encoding $h_{t}$.These encodings are then input into a GNN $G(\cdot)$ to capture the dynamic interdependencies among entities. The output of the GNN $G(\cdot)$ , represented as $y$, encapsulates the spatio-temporal conditions of the entities. Subsequently, $y$ is provided as input to a soft contrastive learning head $H(\cdot)$ equipped with an MLP, which processes it to produce feature representations $c$ in a transformed feature space where similar samples are pulled closer and dissimilar samples are pushed apart. Finally, these transformed spatio-temporal conditions $c$ are input into an entity-aware normalization flow model $F(\cdot)$ for density estimation. (b) The aim of the proposed soft contrastive learning(SCL) head, is mapping the semantically similar nodes together.
@@ IV-A Notation and Problem Definition
IV-B Neural Network Structure of USD
The core idea of USD is to learn more accurate latent spatial representations using a combination of data augmentation and soft contrast learning. Fig. 2 illustrates the details of USD framework, which synergizes RNN $R(\cdot;\theta)$, GNN $G(\cdot;\phi)$, and FLOW model $F(\cdot;\alpha)$ to adeptly navigate the complexities of multivariate time series data for fault detection. In summary, the RNN $R(\cdot;\theta)$ captures the temporal correlations within the data, generating time encodings that are then input into the GNN $G(\cdot;\phi)$ to capture the dynamic interdependencies among entities. The output of the GNN $G(\cdot;\phi)$, representing the spatio-temporal conditions of the entities, is fed into the entity-aware normalizing flow model $F(\cdot;\alpha)$ for density estimation, enabling the model to discern normal and abnormal states with precision. The specific detailed definitions are as follows:
Temporal Dynamics with RNN [$R(\cdot;\theta)$]. We model the temporal variations of each entity using an RNN $R(\cdot)$. The hidden state $g^{c}$ at time window $c$ is updated as
$g^{c}={R}\left(x^{c};\theta\right),$ (3)
where $x^{c}$ represents the input features at time window $c$, the $\theta$ is the parameter of RNN, and $g^{c}$ is the hidden state of the RNN at time window $c$.
Dynamic Interdependencies with GNN [$G(\cdot;\phi)$]. A graph structure learning module captures the dynamic interdependencies among entities. The graph convolution operation, applied to the time encoding from the RNN and the learned graph structure $G$, is given by
$y^{c}={G}\left(g^{c};\phi\right),$ (4)
where $y^{c}$ represents the spatio-temporal condition of the entities at time window $c$, $\phi$ is the parameter of GNN, and $y^{c}$ is the output of the GNN at time window $c$. The graph structure learning module captures the evolving interdependencies among entities, enabling the model to discern the complex relationships and interactions that underpin the multivariate time series data.
Soft Contrastive Learning Head with MLP [$H(\cdot;\omega)$]. To refine the feature representations and enhance the discriminative power of the model, a Soft Contrastive Learning Head implemented with a Multilayer Perceptron (MLP) is employed. This component $H(\cdot)$ takes the output $y^{c}$ from the GNN as input and processes it to generate a transformed feature space where similar samples are brought closer, and dissimilar samples are pushed apart. The transformation is defined as:
$z^{c}=H(y^{c};\omega),$ (5)
where $z^{c}$ represents the contrastively learned feature representation of the entities at time window $c$, and $\omega$ denotes the parameters of the MLP. The objective of this component is to maximize the margin between different classes in the feature space, effectively improving the separability and robustness of the model. This is achieved through a contrastive loss function, which penalizes similar pairs that are too far apart and dissimilar pairs that are too close, thereby fine-tuning the embeddings to better capture the nuances of the underlying data dynamics.
Density Estimation with FLOW Model [$F(\cdot;\alpha)$]. The spatio-temporal conditions are then input into the entity-aware normalizing flow model for fine-grained density estimation. The transformation of a data point $z$ into $\hat{z}$ through the flow model is described as
$\hat{z}=f(z;\alpha),$ (6)
where $f$ represents the flow transformation parameterized by $\alpha$.
IV-C Maximum Likelihood Estimation (MLE) Loss
To enhance the anomaly detection capabilities of the USD framework, which includes the RNN, GNN, and FLOW models, the entire system is optimized using Maximum Likelihood Estimation (MLE). This optimization approach ensures the models are finely tuned to detect anomalies effectively:
$\mathcal{L}_{\text{MLE}}=\frac{1}{N}\sum_{c=1}^{N}-\frac{1}{2}\left\|\hat{z}^{c}-\mu\right\|_{2}^{2}+\log\left|\operatorname{det}\frac{\partial f_{\alpha}}{\partial x^{c}}\right|,$ (7)
where $N$ represents the total number of time windows, $K$ denotes the number of entities, and $\mu_{k}$ is the mean vector of the target distribution $\mathcal{Z}$. The MLE loss function is crafted to maximize the probability of the transformed data points $\hat{z}_{i}$ being observed under the model’s assumptions, thereby ensuring that the model faithfully represents the true distribution of the data.
IV-D Data Augmentation for Soft Contrastive Learning
@@
Data Augmentation by Linear Interpolation. Data augmentation strategy augments the dataset by creating intermediate samples between existing data points [28]. For two points $x^{c}$ and $x^{c}_{\text{new}}$, a new sample $x_{\text{new}}^{c}$ is formulated as:
@@
Together, neighborhood discovery and linear interpolation form a powerful data augmentation framework that leverages both the local structure and the linear relationships within the data. This approach not only expands the dataset but also ensures that the augmented samples are meaningful and representative of the original data characteristics, thereby improving the model’s performance and robustness.
IV-E Soft Contrastive Learning Loss with Data Augmentation
In data-augmentation-based contrastive learning (CL), the task is framed as a binary classification problem over pairs of samples. Positive pairs, drawn from the joint distribution $(x^{\text{c1}},x^{c2})\sim P_{x^{\text{c1}},x^{c2}}$, are labeled as $\mathcal{H}_{c1,c2}=1$, whereas negative pairs, drawn from the product of marginals $(x^{\text{c1}},x^{c2})\sim P_{x^{\text{c1}}}P_{x^{c2}}$, are labeled as $\mathcal{H}_{ck}=0$. The goal of contrastive learning is to learn representations that maximize the similarity between positive pairs and minimize it between negative pairs, utilizing the InfoNCE loss.
$\mathcal{L}_{\text{CL}}\left(x^{\text{c1}},x^{c2},\{x^{\text{cn}}\}_{cn=1}^{N_{K}}\right)$ $\!=\!-\!\log\!\frac{\exp({{z^{\text{c1}}}^{T}\!z^{\text{c2}}})}{\sum_{k=1}^{N_{K}}\!\exp({{z^{\text{c1}}}^{T}\!z^{\text{cn}}})}$ (10)
$=\!-\!\log\!\frac{\exp(S(z^{\text{c1}},\!z^{\text{c2}}))}{\sum_{k=1}^{N_{K}}\exp(S(z^{\text{c1}},\!z^{\text{cn}}))},$

### v2

# Deep Multi-Manifold Transformation Based Multivariate Time Series Fault Detection

@@ Abstract
Unsupervised fault detection in multivariate time series plays a vital role in ensuring the stable operation of complex systems. Traditional methods often assume that normal data follow a single Gaussian distribution and identify anomalies as deviations from this distribution. However, this simplified assumption fails to capture the diversity and structural complexity of real-world time series, which can lead to misjudgments and reduced detection performance in practical applications. To address this issue, we propose a new method that combines a neighborhood-driven data augmentation strategy with a multi-manifold representation learning framework. By incorporating information from local neighborhoods, the augmentation module can simulate contextual variations of normal data, enhancing the model’s adaptability to distributional changes. In addition, we design a structure-aware feature learning approach that encourages natural clustering of similar patterns in the feature space while maintaining sufficient distinction between different operational states. Extensive experiments on several public benchmark datasets demonstrate that our method achieves superior performance in terms of both accuracy and robustness, showing strong potential for generalization and real-world deployment.
Index Terms:
Unsupervised Soft Contrastive Learning, Fault Detection, Multivariate Time Series, Data Augmentation
@@ I Introduction
Unsupervised fault detection [23, 15] in multivariate time series has become increasingly vital in both academic research and practical domains, especially within industrial environments and the management of large-scale systems [29, 21, 36]. By eliminating the need for labeled data, it enables early identification of anomalies and incipient faults [38, 6], which is crucial for minimizing maintenance overhead and averting potential system failures [17]. This capability makes it an indispensable tool for real-time monitoring and predictive maintenance in scenarios where labeled samples are scarce or entirely unavailable [40].
Figure 1: Motivation of DMTFD. (a) Existing methods rely on Gaussian assumptions and fail to capture complex latent structures with multiple sub-states. (b) Our proposed DMTFD introduces multi-manifold assumption to better model state variations and improve fault detection accuracy.
To address this challenge, a wide range of approaches have been explored, from classical statistical methods to modern deep learning-based models [37, 32]. Despite the notable progress made in recent years [20], most existing techniques still exhibit unsatisfactory detection performance in practical applications, characterized by frequent false positives and missed anomalies (see Table I and Table II). This is largely due to the restrictive assumptions imposed during modeling, often necessitated by the absence of supervision. A prevailing strategy is to model normal behavior as a Gaussian distribution and to flag deviations as anomalies [39], as illustrated in Fig. 1(a). While computationally convenient, such assumptions fail to capture the intrinsic diversity and nuanced dynamics of real-world systems.
In reality, both normal and anomalous behaviors often comprise multiple heterogeneous sub-patterns, each exhibiting distinct temporal and structural traits (Fig.1(b)). Relying on a single Gaussian model obscures these internal variations, leading to suboptimal representations and a reduced capacity to detect subtle deviations [23]. This oversimplification not only impairs the model’s ability to capture gradual state transitions but also undermines robustness in highly variable environments[10, 25]. Consequently, there is a pressing need to move beyond Gaussian-centric assumptions and adopt more expressive modeling paradigms capable of characterizing complex, multi-modal system behaviors.
To address the limitations caused by the oversimplified Gaussian distribution assumption, we propose an unsupervised framework, termed DMTFD (Deep Multi-manifold Transformation for Fault Detection). Unlike prior methods that characterize normal states with a single Gaussian distribution [39], DMTFD adopts a multi-manifold assumption, which acknowledges the intrinsic diversity of normal operating conditions across different system states (see Fig. 1(b)).
Concretely, DMTFD leverages a multi-manifold transformation module that maps multivariate time series into a latent space where heterogeneous local patterns are better disentangled and aligned. This transformation enables the model to preserve fine-grained distinctions between sub-manifolds of normal behaviors while amplifying their deviation from anomalous ones. To construct a meaningful representation space under this assumption, we introduce a neighbor-aware augmentation strategy that generates context-consistent variants of each sample [18], and adopt a softened similarity constraint to gently pull semantically close instances. This design alleviates the rigidity of binary contrastive signals and supports more nuanced anomaly boundary modeling [41].
Comprehensive experiments show that DMTFD significantly outperforms existing methods in both AUC and PR metrics. On benchmark datasets, DMTFD achieves consistently lower false alarm rates and higher detection accuracy. Furthermore, our visualization analysis supports the plausibility of the multi-Gaussian assumption and illustrates how soft contrastive learning effectively captures underlying sub-state structures. (a) Multi-manifold modeling of operational states. We break away from the conventional single-Gaussian assumption and adopt a multi-manifold view that reflects the heterogeneity of both normal and abnormal conditions. This insight leads to a more faithful and flexible modeling of system behaviors. (b) New framework for unsupervised representation. We propose a novel framework that integrates neighbor-based data augmentation and loss function to construct a smooth and discriminative latent space. (c) Superior empirical performance across benchmarks. Our method achieves over 5% improvement in detection metrics compared to existing baselines, demonstrating strong robustness and generalization in real-world industrial scenarios.
@@ II-A Time Series Anomaly Detection
In time series anomaly detection, One-Class Classification (OCC) methods—such as USAD [4], and DAEMON [8]—typically assume access to purely normal data during training [5, 16]. GANF [9] employs graph neural networks for anomaly detection in MTS. MTGFlow [39] combines dynamic graphs with normalizing flows, while MTGFlow-cluster [40] further boosts accuracy by clustering entities. AnomalyLLM [22] leverages large language models for knowledge distillation, using prototypical signals and synthetic anomalies to achieve a improvement on UCR datasets across 15 benchmarks. PeFAD [31] introduces PLM-based parameter-efficient federated learning with anomaly masking and synthetic distillation, improving performance by up to 28.74%. Graph-MoE [15] enhances GNN-based detection via expert mixtures and memory routers, effectively utilizing hierarchical features. In addition, prior work USCL [23] proposes a contrastive learning framework tailored for multivariate time series fault detection, addressing limitations of hard contrastive loss in the presence of view-level noise. Building upon this, the present work introduces a multi-manifold transformation strategy and a generalized similarity modeling approach, enabling more expressive representations and significantly improving detection performance and robustness across benchmarks.
@@ II-B Contrastive Learning and Soft Contrastive Learning
Soft contrastive learning [35, 33] and deep manifold learning [24, 19] represent two advanced methodologies in the domain of machine learning, particularly in the tasks of unsupervised learning and representation learning. These techniques aim to leverage the intrinsic structure of the data to learn meaningful and discriminative features. At its core, soft contrastive learning [35, 33, 28] is an extension of the contrastive learning framework [7, 13, 14], which aims to learn representations by bringing similar samples closer and pushing dissimilar samples apart in the representation space. However, soft contrastive learning introduces a more nuanced approach by incorporating the degrees of similarity between samples into the learning process, rather than treating similarity as a binary concept. This is achieved through the use of soft labels or continuous similarity scores, allowing the model to learn richer and more flexible representations. The softness in the approach accounts for the varying degrees of relevance or similarity among data points, making it particularly useful in tasks where the relationship between samples is not strictly binary or categorical, such as in semi-supervised learning or in scenarios with noisy labels.
@@ III Preliminary
Figure 2: Overview of the proposed DMTFD framework. (a) Temporal sequences $x,x^{+}$ are processed via RNN to capture temporal dynamics, followed by a GNN to model inter-entity dependencies, yielding embeddings $y$. A soft contrastive learning head maps $y$ to transformed features $c$, which are then used for density modeling via a flow-based generative network. (b) The SCL head helps cluster similar samples and separate dissimilar ones in the feature space.
@@ IV-A Notation and Problem Definition
IV-B Neural Network Structure of DMTFD
The core idea of DMTFD is to learn precise and discriminative latent representations by integrating data augmentation with soft contrastive learning. As illustrated in Fig. 2, the DMTFD framework jointly leverages a Recurrent Neural Network (RNN) $R(\cdot;\theta)$, a Graph Neural Network (GNN) $G(\cdot;\phi)$, and a normalizing flow model $F(\cdot;\alpha)$ to effectively model the temporal patterns, relational structures, and distributional properties of multivariate time series data for fault detection.
RNN-GNN-Flow Architecture. To jointly capture temporal dynamics, inter-entity dependencies, and density patterns, we adopt an RNN-GNN-Flow architecture. Given the multivariate input at time window $c$, denoted as $x^{c}$, we first compute the temporal encoding $g^{c}=R(x^{c};\theta)$ using a recurrent neural network parameterized by $\theta$; a graph neural network with parameters $\phi$ then produces the spatio-temporal embedding $y^{c}=G(g^{c};\phi)$; finally, a normalizing flow with parameters $\alpha$ estimates the log-likelihood score $\ell^{c}=F(y^{c};\alpha)$, which is used to detect anomalies—where low likelihood indicates high abnormality.
Manti-manifold Transformation Head (MTH). To enhance the discriminative quality of the learned representations, we introduce a Multi-manifold Transformation Head, implemented as a multilayer perceptron (MLP) $H(\cdot;\omega)$. This module receives the spatio-temporal embedding $y^{c}$ from the GNN as input and projects it into a latent feature space optimized for contrastive learning: $z^{c}=H(y^{c};\omega),$ where $\omega$ denotes the learnable parameters of the MLP, and $z^{c}$ is the contrastive representation corresponding to time window $c$. The goal of this module is to enforce semantic alignment between similar instances while ensuring separation among dissimilar ones. This is achieved via a soft contrastive loss, which penalizes embeddings of positive pairs (semantically similar samples) that are far apart and encourages larger distances for negative pairs (semantically dissimilar samples) that are too close. Unlike hard contrastive formulations that impose strict binary similarity, the soft version allows for graded similarity relationships, thereby providing a smoother optimization landscape and improving robustness to intra-class variability.
IV-C Maximum Likelihood Estimation (MLE) Loss and Data Augmentation
A key objective of the DMTFD framework is to accurately identify abnormal behaviors by modeling the underlying distribution of normal operational data. To this end, we employ Maximum Likelihood Estimation (MLE) as the training criterion for the flow-based density estimation module. The core motivation behind this design is to encourage the model to assign high likelihoods to normal patterns while assigning low likelihoods to anomalous deviations, thus enabling precise and fine-grained fault detection.
Formally, given the contrastive representation $z^{c}$ for time window $c$, the normalizing flow transformation $f(\cdot;\alpha)$ maps it into a latent variable $\hat{z}^{c}=f(z^{c};\alpha)$ that follows a known target distribution (typically standard Gaussian with mean $\mu$ and identity covariance). The MLE loss is then defined as:
$\mathcal{L}_{\text{MLE}}=\frac{1}{N}\sum_{c=1}^{N}\left[-\frac{1}{2}\left\|\hat{z}^{c}-\mu\right\|_{2}^{2}+\log\left|\det\left(\frac{\partial f_{\alpha}}{\partial z^{c}}\right)\right|\right],$ (3)
where $N$ is the number of time windows, and $\mu$ is the mean of the target Gaussian distribution $\mathcal{Z}$. The first term encourages $\hat{z}^{c}$ to match the target distribution, while the second term—i.e., the log-determinant of the Jacobian—ensures that the flow transformation remains invertible and probability-preserving. By optimizing $\mathcal{L}_{\text{MLE}}$, the flow model learns to model the density landscape of normal patterns precisely. At inference time, data points with significantly low log-likelihoods under the learned distribution are flagged as anomalies, enabling accurate and interpretable fault detection across diverse temporal and relational contexts.
@@ IV-C Maximum Likelihood Estimation (MLE) Loss and Data Augmentation
Data Augmentation by Linear Interpolation. Data augmentation strategy augments the dataset by creating intermediate samples between existing data points. For two points $x^{c}$ and $x^{c}_{\text{new}}$, a new sample $x_{\text{new}}^{c}$ is formulated as,
@@
IV-D Multi-Manifold Loss with Data Augmentation
In data-augmentation-based contrastive learning (CL), the task is framed as a binary classification problem over pairs of samples. Positive pairs, drawn from the joint distribution $(x^{\text{c1}},x^{c2})\sim P_{x^{\text{c1}},x^{c2}}$, are labeled as $\mathcal{H}_{c1,c2}=1$, whereas negative pairs, drawn from the product of marginals $(x^{\text{c1}},x^{c2})\sim P_{x^{\text{c1}}}P_{x^{c2}}$, are labeled as $\mathcal{H}_{ck}=0$. The goal of contrastive learning is to learn representations that maximize the similarity between positive pairs and minimize it between negative pairs, utilizing the InfoNCE loss. The $\mathcal{L}_{\text{CL}}\left(x^{\text{c1}},x^{c2},\{x^{\text{cn}}\}_{cn=1}^{N_{K}}\right)=$
$-\log\!\frac{\exp({{z^{\text{c1}}}^{T}\!z^{\text{c2}}})}{\sum_{k=1}^{N_{K}}\!\exp({{z^{\text{c1}}}^{T}\!z^{\text{cn}}})}=\!-\!\log\!\frac{\exp(S(z^{\text{c1}},\!z^{\text{c2}}))}{\sum_{k=1}^{N_{K}}\exp(S(z^{\text{c1}},\!z^{\text{cn}}))},$ (6)